# GMValuator: Similarity-based Data Valuation for Generative Models

### v7

@@ IV-B Image Quality Assessment
1:  // Generate the synthetic dataset
2:  $\hat{X}=\{\hat{x}_{j}\}_{j=1}^{m}\leftarrow\mathcal{A^{*}}(z_{j})$, for $z_{j}\in\mathcal{Z}$
3:  for $\hat{x}_{j}$ in $\hat{X}$ do
4:     // Matching process (see Sec. IV-A)
5:     $\mathcal{P}_{j}=f(X,\hat{x}_{j})$  // Including two phases
6:     for $x_{i}$ in $\mathcal{P}_{j}$ do
7:        $d_{ij}\leftarrow{\rm LPIPS}(x_{i},\hat{x}_{j})$ or ${\rm DreamSim}(x_{i},\hat{x}_{j})$
8:     end for
9:     // Image Quality Assessment (see Sec. IV-B)
10:     $q_{j}=MANIQA(\hat{x}_{j})$
11:     // Contribution Score Calculation (see Sec. IV-C)
12:     Calculate score $\mathcal{V}(x_{i},\hat{x}_{j},d_{ij},q_{j})$ using Eq. (4)
13:  end for
14:  // Calculation of data value and return the result $\Phi$
15:  for $x_{i}$ in $X$ do
16:     Calculate $x_{i}$’s value $\phi_{i}$ using Eq.(3)
17:  end for
18:  return $\Phi=\left\{\phi_{1},\phi_{2},...,\phi_{n}\right\}$
IV-C Value Calculation
Finally, we utilize the image quality $q$ to calibrate the contribution scores. Notably, during the recall phase of similarity matching, we select the top $k$ contributors $\mathcal{P}_{j}$ for the generated data $\hat{x}_{j}$. Consequently, we only consider the contribution score between $\mathcal{P}_{j}$ and $\hat{x}_{j}$ by setting $\mathcal{V}(x_{i},\hat{x}_{j})=0$ for $x_{i}\notin\mathcal{P}_{j}$. This strategy addresses challenge 3 by assigning zero scores to irrelevant samples, effectively reducing bias and noise in value estimation when dealing with a large $n$. Hence, we define the contribution score function of each training data point for a specific synthetic data point as follows:
which is used to give different scores to training data points according to their ranking distance. The final data value is obtained by plugging Eq. (4) into Eq. (3).
V Experiments
  • C1: Identical Class Test. Following the concept of an identical class [30], we posit that the most significant contributors among the training samples should belong to the same class as the generated sample produced by a well-trained generative model, which is also used for evaluation in VAE-TrancIn [14].
  • C2: Identical Attributes Test. We consider the attributes of samples as the ground truth to examine the most significant contributors by our approach. We examine the overlap level of attributes (i.e., the number of identical attributes) between the most valuable training samples with the generated sample
  • C3: Data Cleansing. Under the assumption, noisy training data samples can be identified as low-value training data. To evaluate the performance of data valuation approaches, we consider the value level of noisy data samples as the performance metric across various approaches.
TABLE I: Experiment settings.
Dataset Network
C1 MNIST, CIFAR-10 GAN, Diffusion, $\beta$-VAE
C2 CelebA Diffusion-StyleGAN
AFHQ StyleGAN
C3 Noise MNIST DCGAN
C4: Hardware Evironment
CPU One RTX 3080 (10GB) GPU
GPU
12 vCPU Intel(R) Xeon(R),
Platinum 8255C CPU @ 2.50GHz
Experiment Setup. Our experiment settings are listed in in Table I. Particularly, we consider different type of generative models, including GAN based models, $\beta$-VAE [31], and diffusion models [32, 33]. The generation tasks are conducted on benchmark datasets (i.e., MNIST [34] and CIFAR [35]), face recognition dataset (i.e., CelebA [36]) and high-resolution image dataset with size 512 $\times$ 512, and 1024 $\times$ 1024 (i.e., AFHQ [37], FFHQ [38]).
Note that GMValuator is a model agnostic method. To showcase its efficacy, we compare it with two baseline methods: VAE-TracIn [14] and IF4GAN [12]. VAE-TracIn identifies the most significant contributors for a generated sample, which is same to the similarity matching process of GMValuator (Sec. IV-A). Thus, we compare GMValuator with VAE-TracIn in Sec. V-A and Sec. V-B. IF4GAN finds the high-valued training samples for generative model training, which we regard it as baseline method in section V-C. We also examine our approach using DreamSim, LPIPS and $l_{2}$-distance in the re-ranking phase, referred to as GMValuator (DreamSim), GMValuator (LPIPS) and GMValuator ($l_{2}$-distance) respectively, while the GMValuator without the re-ranking phase is referred to as GMValuator (No-Rerank).
V-A Identical Class Test (C1)
Methodology. In this subsection, we follow setup of VAE-TracIn [14] by training separate $\beta$-VAE models on MNIST [35], CIFAR [34]. We then attribute the most significant contributors by VAE-TracIn and our approach over generated samples. By the concept of identical class test, we expects a perfect data valuation approach should indicate training data in the same subclass of the generated sample $\hat{x}_{j}$ contribute more to $\hat{x}_{j}$. We examine GMValuator (DreamSim), GMValuator (LPIPS), GMValuator ($l_{2}$-distance) and GMValuator (No-Rerank) separately. Apart from that, due to our method is model-agnostic, we also valid our method on GAN, Diffusion model, which are other kinds of generative models.
Results. For a given generated data, we examine the class(es) of is top $k$ contributors in the training data. We count the number of training $Q$ samples in the top $k$ contributors, which has the identical class as the generated data. The identical class ratio $\rho=Q/k$. We report the averaged $\rho$ over the generated datasets (the data size $m$=100) on different choices of $k$ in Table  II. GMValuator has the most contributors that belong to the same class with the generated sample among on MNIST and CIFAR-10, while GMValuator (DreamSim) has the best performance. The results for CIFAR-10 using VAE-TracIn is extremely bad. This could be attributed to underfitting, as the authors mentioned in their study [14], training a good $\beta$-VAE model on CIFAR-10 is challenging. The results of other diverse generative models are show in Table VI in appendix.
TABLE II: Performance comparison of Identical Class Test.
MNIST (%) $k$=30 $k$=50 $k$=100
VAE-TracIn 72.00 71.11 68.58
GMValuator (No-Rerank) 86.41 85.13 85.95
GMValuator ($l_{2}$-distance) 87.76 86.95 85.92
GMValuator (LPIPS) 88.78 88.19 86.69
GMValuator (DreamSim) 88.78 88.05 86.84
CIFAR-10 (%) $k$=30 $k$=50 $k$=100
VAE-TracIn 6.28 3.77 1.88
GMValuator (No-Rerank) 72.66 72.25 70.71
GMValuator ($l_{2}$-distance) 72.66 72.25 70.71
GMValuator (LPIPS) 72.60 71.75 70.84
GMValuator (DreamSim) 77.94 76.41 73.47
Figure 4: Visualization for Identical Attributes Test on CelebA. Left: generated samples. Right: top $k$ contributors.
V-B Identical Attributes Test (C2)
Methodology. By regarding some attributes of images as the ground-truth, we assume that the most significant contributors should have the similar attributes to the generated sample. We train Diffusion-StyleGAN [33] on CelebA, and leverage our approach to identify the most significant contributors in training samples, and use some some attributes including glasses, hair, hat, sex as the ground-truth, to check the correctness of our approach for identifying the most significant contributors for a generated sample. We also examine this on AFHQ, FFHQ dataset by StyleGAN [37].
Results. The results in Table III indicates that GMValuator can find the most significant contributors with the same attributes to the generated sample. Besides, using DreamSim in similarity re-ranking phase obtains the best performance over these three attributes. We also visualize some visible results in Figure 4 on CelebA and Figure 6 on AFHQ, FFHQ in Appendix. As we can see, the most significant contributors have the similar attributes (e.g., skin color, hair) to the generated sample.
TABLE III: Performance of Identical Attributes Test of some attributes including Hat, Sex, and Eyeglasses on CelebA.
Top K contributors: $k$=5 $k$=10 $k$=15
Attribute: Hat (%)
GMValuator (No-Rerank) 92.52 92.12 91.92
GMValuator ($l_{2}$-distance) 90.91 89.90 89.83
GMValuator (LPIPS) 96.77 96.57 96.90
GMValuator (DreamSim) 97.78 97.07 96.90
Attribute: Gender (%)
GMValuator (No-Rerank) 75.15 75.25 75.82
GMValuator ($l_{2}$-distance) 63.64 61.01 60.00
GMValuator (LPIPS) 97.98 97.48 97.37
GMValuator (DreamSim) 99.19 98.99 98.79
Attribute: Eyeglasses (%)
GMValuator (No-Rerank) 91.52 91.52 91.45
GMValuator ($l_{2}$-distance) 94.95 94.65 93.80
GMValuator (LPIPS) 94.95 94.65 93.80
GMValuator (DreamSim) 96.77 96.26 95.96
V-C Data Cleaning (C3)
Methodology In this experiment, we introduce 100 noisy images into the MNIST dataset to create a new corrupted training dataset and each noisy image is generated by adding Gaussian noise to the clean data. These contaminated samples can diminish the performance of generative models, and as a result, they are anticipated to possess lower values compared to the rest of the training dataset. Our objective is to assess the performance of both GMValuator and its baseline method, IF4GAN [12], by analyzing the value rankings of noisy samples. An effective approach should place noisy data in lower positions within the value ranking. We conducted this evaluation using 10,000 generated images and referred to the study [12] for the baseline method (IF4GAN).
Results. The results are depicted in Figure 5 demonstrate that the values of all noisy training samples, as calculated by GMValuator, are lower compared to the values calculated by IF4GAN. This observation suggests that the performance of GMValuator is significantly better than that of IF4GAN. Besides, the performance of GMValuator with re-ranking phase is better than GMValuator (No-rerank).
V-D Efficiency (C4)
Methodology. We thoroughly evaluate and compare the efficiency of our approach against existing data valuation methods for generative models discussed above: VAE-TracIn and IF4GAN. Specifically, we measure the attribute time for one generated sample on an average of 100 test samples using both VAE-TracIn and our proposed approach on MNIST and CIFAR-10. Additionally, under the same setting in C3, we assess the data valuation time when utilizing IF4GAN and our approaches on noise MNIST which we mentioned in C3.
Results. The average time taken to attribute the most significant contributors for one generated sample is calculated and compared with the baseline method (VAE-TracIn), as demonstrated in Table IV. Our approaches are all greatly more efficient than the baseline methods on both datasets. When it comes to data valuation for GAN, GMValuator (No-rerank) and GMValuator (LPIPS) are significantly better than IF4GAN. The aforementioned phenomenon is attributed to the costly nature of the Hessian estimation process, despite the utilization of certain acceleration methods. It is noticable that GMValuator (DreamSim) is lightly time-consuming when compare to IF4GAN. This lead to a trade off problem due to the GMValuator using DreamSim in re-ranking phase obtained the better performance from C1 to C3.
Overall, through the above experiments, GMValuator outperform than baseline methods, while GMValuator (DreamSim) obtains the best performance. This is due to the fact that DreamSim capture mid-level similarities in image semantic content and layout compare to LPIPS. And the selection of different GMValuators can be combined with the consideration of their efficiency.
TABLE IV: Efficiency Comparison
Attribute for VAE Time(s)
MNIST VAE-TracIn 47.945
GMValuator (No-Rerank) 0.250
GMValuator ($l_{2}$-distance) 0.339
GMValuator (LPIPS) 0.477
GMValuator (DreamSim) 1.709
CIFAR-10 VAE-TracIn 66.178
GMValuator (No-Rerank) 0.755
GMValuator ($l_{2}$-distance) 1.226
GMValuator (LPIPS) 2.412
GMValuator (DreamSim) 15.491
Data Valuation for GAN Time(s)
Noise MNIST IF4GAN 14,543
GMValuator (No-Rerank) 2,137
GMValuator ($l_{2}$-distance) 3,388
GMValuator (LPIPS) 4,771
GMValuator (DreamSim) 17,086
VI Conclusion
Appendix A Statistical Results for Figure 1
It is evident by visualization in Figrue 1 that the data points in $X_{v2}$ (used for training) are more overlapped with generated data than data points in $X_{v1}$ (not used for training). We perform statistic testing on data values obtained by GMvaluator, to examine if data points $X_{v2}$ (used for training) have significant higher values than those of the data points in $X_{v1}$ (not used for training).
TABLE V: The statistic test of data values of $X_{v1}$ versus $X_{v2}$ using different generative models. $X_{v2}$ is supposed to have higher value than $X_{v1}$, given the generated data.
@@ Appendix A Statistical Results for Figure 1
BigGAN
Average value (v1) 0.319654
Average value (v2) 1.632352
P-value 6.937027 ×$\times 10^{-68}$
T-statistic 17.924512
Significance level 0.01
@@
Classifier-free Guidance Diffusion
Average value (v1) 0.030434
Average value (v2) 0.369565
P-value 8.053195 ×$\times 10^{-55}$
T-statistic 15.947860
Significance level 0.01
Result
p-value less than 0.01, reject $H_{0}$
To this end, we use a t-test with null hypothesis that data values in $X_{v1}$ should not be samller than those of $X_{v12}$. We compute a $p$-value, which is the probability of getting a difference as large as we observed, or larger, under the null hypothesis. If the $p$-value is very low, we reject the null hypothesis and consider our approach, GMValuator, to be verified with a high level of confidence (1-$p$). Typically, a $p$-value smaller than significance level 0.01 is used as a threshold for rejecting the null hypothesis. Table V showcases the outcomes of $X_{v1}$ and $X_{v2}$ in CIFAR-10 with $p\ll 0.01$ for both BigGAN and diffusion model, indicating that the data points in $X_{v2}$ have significantly more value than those in $X_{v1}$. Consequently, these findings align with the presumption that the trained dataset $X_{v2}$ has a higher value than the untrained dataset $X_{v1}$ and verify our approach.
Figure 6: Visualization of Identical Attributes Test on AFHQ and FFHQ. The results shown in the first and second subfigures on the left are conducted on AFHQ-Cat and AFHQ-Dog, respectively. The subfigure on the right presents the results on FFHQ. In each subfigure, the generated samples are on the left, and the top $k$ contributors in the training dataset are on the right.
Appendix B Additional Results on C1 and C2
TABLE VI: Performance comparison of Identical Class Test.
@@ Appendix B Additional Results on C1 and C2
B-A Identical Class Test on Other Generative Models
We have presented Indetical Class Test (C1) on $\beta$-VAE and MNIST [34], CIFAR-10 [35] in Sec V-A in our main context. Since GMValuator is model-agnostic, we further validate our method of C1 on other generative models.
Here, we conduct the experiments using a GAN and a Diffusion Model on MNIST. The architectural details of the used generative models are described in Sec E-B in Appendix. We also conduct experiment on BigGAN [25] and Class-free Guidance Diffusion [32] with CIFAR-10. We used the same number of generated samples $m=100$ as the experiments presented in Sec V.
Following the similar settings in Sec. V-A (C1), we examine the class(es) of is top $k$ contributors for a given generated data in the training data. We calculate the number of training samples, denoted as $Q$, from the top $k$ contributors that have the same class as the generated data. The identical calss ratio, denoted as $\rho$, is calculated as $\rho=Q/k$. We report the average value of $\rho$ across the generated datasets for different choices of $k$ in Table VI. GMValuator (DreamSim) has the highest ratio of contributors that belong to the same class as the generated sample among most of the models evaluated on MNIST and CIFAR-10 datasets for different values of $k$. And the ratio improves as the value of $k$ decreases, which is consistent to the top $k$ assumption and validates our method.
B-B Identical Attributes Test on Other Datasets
To further validate GMValuator, we also conducted experiments on high-resolution datasets: AFHQ [37] and FFHQ [38], following the same settings in Sec V-B. The results are shown in Figure 6, which demonstrates the effectiveness of our methods in C2. The results show that the top $k$ contributors have the similar attributes with the generated sample such as the fur color of cats or dogs in AFHQ. For the experiment conducted on FFHQ, human faces attributes of the most significant contributors are also similar to the attributes in generated images.
@@ B-B Identical Attributes Test on Other Datasets
Appendix C Different Generated Data Sizes

### v8

@@ Appendix B Algorithm
1:  $\hat{X}=\{\hat{x}_{j}\}_{j=1}^{m}\leftarrow G^{*}(z_{j})$, for $z_{j}\in\mathcal{Z}$  // Generate the synthetic dataset
2:  for $\hat{x}_{j}$ in $\hat{X}$ do
3:     // Matching process (see Sec. 3.1)
4:     $\mathcal{P}_{j}=f(X,\hat{x}_{j})$  // Including two phases
5:     for $x_{i}$ in $\mathcal{P}_{j}$ do
6:        $d_{ij}\leftarrow$ ${\rm DreamSim}(x_{i},\hat{x}_{j})$ or others
7:     end for
8:     $q_{j}=MANIQA(\hat{x}_{j})$  // Image Quality Assessment (see Sec. 3.2)
9:     Calculate score $\mathcal{V}(x_{i},\hat{x}_{j},d_{ij},q_{j})$ using Eq. equation 6 // Contribution Score Calculation (see Sec. 3.3)
10:  end for
11:  // Calculation of data value and return the result $\Phi$
12:  for $x_{i}$ in $X$ do
13:     Calculate $x_{i}$’s value $\phi_{i}$ using Eq.equation 5
14:  end for
15:  return $\Phi=\left\{\phi_{1},\phi_{2},...,\phi_{n}\right\}$
Appendix C Related Work
C.1 Data Valuation
There are three lines of methods on data valuation: metric-based methods, influence-based methods and data-driven methods. In terms of metric-based methods, the commonly-used approach is to calculate its marginal contribution (MC) based on performance metrics (e.g., accuracy, loss). As the basic method depending on performance metrics for data valuation, LOO (Leave-One-Out) Cook (1977) is used to evaluate the value of the training sample by observational change of model performance when leaving out that data point from the training dataset. To overcome inaccuracy and strict desirability of LOO, SV Ghorbani & Zou (2019) and BI Wang & Jia (2023) originated from Cooperative Game Theory are widely used to measure the contribution of data Jia et al. (2019b); Ghorbani et al. (2020). Considering the joining sequence of each training data point, SV needs to calculate the marginal performance of all possible subsets in which the time complexity is exponential. Despite the introduction of techniques such as Monte-Carlo and gradient-based methods, as well as others proposed in the literature, approximating data significance value (SV) is computationally expensive and it typically requires retraining Ghorbani & Zou (2019); Jia et al. (2019a). The computational cost and need for unconventional performance metrics present difficulties in adapting the methods to generative models. As for influence-based methods, they evaluate the influence of data points on model parameters by computing the inverse Hessian for data valuation Jia et al. (2019a); Richardson et al. (2019); Saunshi et al. (2022). Due to the high computational cost, some approximation methods have also been proposed Pruthi et al. (2020). In addition, the use of influence function for data valuation is not limited to discriminative models, but can also be applied to specific generative models such as GAN and VAE Terashita et al. (2021); Kong & Chaudhuri (2021). When it comes to data-driven methods, most of them are training-free methods that focus on the data itself Xu et al. (2021); Wu et al. (2022); Just et al. (2023). Additionally, the Information-Theoretic Measures method Boopathy et al. (2023) evaluates the inherent generalization difficulty of tasks and shows potential for adaptation to data valuation. However, it is not applicable to our specific setting due to its reliance on performance metrics, training requirements, and a focus on dataset-level trends.
C.2 Generative Model
Generative models are a type of unsupervised learning that can learn data distributions. Recently, there has been significant interest in combining generative models with neural networks to create Deep Generative Models, which are particularly useful for complex, high-dimensional data distributions. They can approximate the likelihood of each observation and generate new synthetic data by incorporating variations. Variational auto-encoders (VAEs) Rezende et al. (2014) optimize the log-likelihood of data by maximizing the evidence lower bound (ELBO), while generative adversarial networks (GANs) Goodfellow et al. (2020); Karras et al. (2020) involves a generator and discriminator that compete with each other, resulting in strong image generation. Recently proposed diffusion models Ho et al. (2020); Rombach et al. (2022) add Gaussian noise to training data and learn to recover the original data. These models use variational inference and have a fixed procedure with a high-dimensional latent space.
Appendix D Statistical Results for Figure 1
It is evident by visualization in Figure 1 that the data points in $X_{v2}$ (used for training) are more overlapped with generated data than data points in $X_{v1}$ (not used for training). We perform statistic testing on data values obtained by GMValuator, to examine if data points $X_{v2}$ (used for training) have significantly higher values than those of the data points in $X_{v1}$ (not used for training).
Table 6: The statistic test of data values of $X_{v1}$ versus $X_{v2}$ using different generative models. $X_{v1}$ is supposed to have higher value than $X_{v2}$, given the generated data.
@@ Appendix D Statistical Results for Figure 1
BigGAN Classifier-free Guidance Diffusion
Average value (v1) 0.319654 0.030434
Average value (v2) 1.632352 0.369565
P-value 6.937027 ×$\times 10^{-68}$ 8.053195 ×$\times 10^{-55}$
T-statistic 17.924512 15.947860
Significance level 0.01 0.01
@@
p-value less than 0.01, reject $H_{0}$,
value of v2 less than v1 averagely
To this end, we use a t-test with the null hypothesis that data values in $X_{v1}$ should not be smaller than those of $X_{v12}$. We compute $p$-value, which is the probability of getting a difference as large as we observed, or larger, under the null hypothesis. If the $p$-value is very low, we reject the null hypothesis and consider our approach, GMValuator, to be verified with a high level of confidence (1-$p$). Typically, a $p$-value smaller than significance level 0.01 is used as a threshold for rejecting the null hypothesis. Table 6 showcases the outcomes of $X_{v1}$ and $X_{v2}$ in CIFAR-10 with $p\ll 0.01$ for both BigGAN and diffusion model, indicating that the data points in $X_{v2}$ have significantly more value than those in $X_{v1}$. Consequently, these findings align with the presumption that the trained dataset $X_{v2}$ has a higher value than the untrained dataset $X_{v1}$ and verify our approach.
Figure 6: Visualization of Identical Attributes Test on AFHQ and FFHQ. The results shown in the first and second subfigures on the left are conducted on AFHQ-Cat and AFHQ-Dog, respectively. The subfigure on the right presents the results of FFHQ. In each subfigure, the generated samples are on the left, and the top $k$ contributors in the training dataset are on the right.
Appendix E Additional Results on C1 and C2
Table 7: Performance comparison of Identical Class Test.
@@ Appendix E Additional Results on C1 and C2
E.1 (C1) Identical Class Test on Other Generative Models
We have presented Identical Class Test (C1) on $\beta$-VAE and MNIST LeCun et al. (1998), CIFAR-10 Krizhevsky et al. (2009) in Sec. 4.3 in our main context. Since GMValuator is model-agnostic, we further validate our method of C1 on other generative models.
Here, we conduct the experiments using a GAN and a Diffusion Model on MNIST. The architectural details of the used generative models are described in Sec. J.3 in the appendix. We also conduct the experiment on BigGAN Brock et al. (2018) and Class-free Guidance Diffusion Ho & Salimans (2022) with CIFAR-10. We used the same number of generated samples $m=100$ as the experiments presented in Sec. 4.
Following the similar settings in Sec. 4.3 (C1), we examine the class(es) of is top $k$ contributors for a given generated data in the training data. We calculate the number of training samples, denoted as $Q$, from the top $k$ contributors that have the same class as the generated data. The identical class ratio, denoted as $\rho$, is calculated as $\rho=Q/k$. We report the average value of $\rho$ across the generated datasets for different choices of $k$ in Table 7. GMValuator (DreamSim) has the highest ratio of contributors that belong to the same class as the generated sample among most of the models evaluated on MNIST and CIFAR-10 datasets for different values of $k$. And the ratio improves as the value of $k$ decreases, which is consistent with the top $k$ assumption and validates our method.
Table 8: Performance of Identical Attributes Test (C2) of multiple combined attributes.
Top K contributors: $k$=5 $k$=10 $k$=15
Attribute: Eyeglasses & Gender (%)
GMValuator (No-Rerank) 50.10 48.48 48.42
GMValuator ($l_{2}$-distance) 59.19 56.67 56.23
GMValuator (LPIPS) 78.18 74.24 73.87
GMValuator (DreamSim) 92.53 90.61 90.30
Attribute: Eyeglasses & Hat (%)
GMValuator (No-Rerank) 78.79 78.38 85.86
GMValuator ($l_{2}$-distance) 84.44 82.93 83.23
GMValuator (LPIPS) 86.87 86.16 86.33
GMValuator (DreamSim) 89.49 87.58 87.41
Attribute: Gender & Hat (%)
GMValuator (No-Rerank) 58.59 57.47 57.71
GMValuator ($l_{2}$-distance) 61.62 60.51 60.40
GMValuator (LPIPS) 63.84 63.23 62.96
GMValuator (DreamSim) 65.25 64.44 64.18
E.2 (C2) Identical Attributes Test on CelebA
We extend C1 to focus on the attributes present in the images, treating them as ground truth rather than class labels, as discussed in Sec. 4.4. Our experiments now incorporate multiple attributes simultaneously, rather than just a single attribute, to verify the performance of GMValuator. For instance, when evaluating a generated image with both a hat and eyeglasses, the most significant contributors identified should also include both of these attributes. Table 8 showcases some outcomes of identical attributes test when regarding multiple combined attributes as ground truth, which validate the effectiveness of our methods.
E.3 (C2) Identical Attributes Test on Other Datasets
@@
To further validate GMValuator, we also conducted experiments on high-resolution datasets: AFHQ Choi et al. (2020) and FFHQ Karras et al. (2019), following the same settings in Sec 4.4. The results are shown in Figure 6, which demonstrates the effectiveness of our methods in C2. The results show that the top $k$ contributors have similar attributes with the generated sample such as fur color of cats or dogs in AFHQ. For the experiment conducted on FFHQ, human faces attributes of the most significant contributors are also similar to the attributes in generated images.
Appendix F Different Generated Data Sizes